OpenStudy (anonymous):

Find the domain. Please show all of your work and express your answer in interval notation.

OpenStudy (anonymous):

\[\sqrt{x-5}\]

OpenStudy (anonymous):

\[x-5\geq 0\] \[x\geq 5\]

OpenStudy (anonymous):

so do you take -5 to each side first?

OpenStudy (turingtest):

first is this\[\sqrt{x-5}\ge0\]and after squaring both sides it becomes what satellite wrote

OpenStudy (anonymous):

ok

Directrix (directrix):

express your answer in interval notation. [5,∞)